### v2

@@ 5.2 Trade War
Next, we consider a trade war scenario: the U.S. imposes an additional 20-percentage-point tariff on all the manufacturing imports, and every other country retaliates against the U.S. by imposing tariffs on U.S. exports with the same magnitude. As before, the trade war is assumed to come as a surprise shock. Results are displayed in Figure A8 in Supplemental Appendix L. Even with retaliatory tariffs from around the world, U.S. real per capita consumption rises for the first 13 years relative to the baseline, but then turns negative. The saving rate is lower than the baseline equilibrium, which follows the same intuition as in the unilateral tariffs. Implications for sectoral expenditure shares and value-added shares remain the same qualitatively, but are smaller in magnitude. The manufacturing value-added share rises by 0.7–0.8 percentage point relative to the baseline, as compared to 0.9–1.3 percentage points in the case of the unilateral tariff. As the retaliatory scenario is more plausible in the real world, our result suggests that the sectoral implications of tariffs on manufacturing will be more marginal in a quantitative sense.
Welfare implications of the trade war are summarized in the fourth and fifth columns of Table 1. We confirm that under the trade war scenario, all countries are worse off compared to the baseline equilibrium. The U.S. welfare loss is 0.12 percent, which is larger than the welfare losses of most European countries and Japan. Canada has the largest welfare loss of 1.08 percent.
@@ 6 Conclusion
The model is used to simulate a 20-percentage-point increase in U.S. manufacturing tariffs on all trading partners beginning in 2001. We find that the relative price effect and the trade protection effect dominate the income effect. Specifically, U.S. manufacturing value-added share rises by 0.9–1.3 percentage points, while the service sector declines by 0.8–1.0 percentage point between 2001 and 2014. Despite these shifts supporting protectionist objectives, the welfare impact is quantitatively minor. The U.S. is better off only by 0.36 percent at the expense of all other countries. In particular, Canada’s welfare falls by 1.3 percent. If the other countries retaliate with 20 percentage-point higher tariffs, all countries are worse off, with the U.S. welfare loss being 0.12 percent.
@@
Hayato Kato [^1] Graduate School of Economics, the University of Osaka; CESifo Research Network. Email: hayato.kato@econ.osaka-u.ac.jp     Kensuke Suzuki [^2] Department of Economics, Clark University; Economic Research Center, Nagoya University. Email: ksuzuki@clarku.edu     Motoaki Takahashi [^3] Graduate School of Economics, the University of Osaka. Email: takahashi.motoaki@gmail.com
@@ Appendix A Proof of Proposition 1
$\qquad\qquad\frac{\theta}{1+\mu_{H}}\frac{d\ln\mu_{H}}{d\tau_{HF}}-\frac{\theta\pi_{HF}w^{\theta}\left[(1+\theta)(w/b_{HF})^{\theta}+A^{\theta}\left(\theta(1+\tau_{HF})+(b_{HF}b_{FH})^{-\theta}\right)\right]}{\Gamma(1+\tau_{HF})},$
@@
$\qquad\qquad\qquad\quad\ \ =\begin{cases}\geq 0&\text{if}\ \tau_{HF}\leq(\theta\pi_{FF})^{-1}+\Gamma^{\prime}\ \text{or}\ \tau_{HF}\leq\tau_{HF}^{**}\\ <0&\text{if}\ \tau_{HF}>(\theta\pi_{FF})^{-1}+\Gamma^{\prime}\ \text{or}\ \tau_{HF}>\tau_{HF}^{**}\end{cases},$
@@
$d\ln P_{F}^{j}=\pi_{FH}d\ln w,\qquad\pi_{FH}=\left(\frac{wb_{FH}/A_{H}}{P_{F}}\right)^{-\theta}=\frac{(A/b_{FH})^{\theta}}{w^{\theta}+(A/b_{FH})^{\theta}}=1-\pi_{FF},$ $\mathclap{\forall j}$
$d\ln P_{H}^{j}=\pi_{HF}d\ln b_{HF}^{j}+\pi_{HH}d\ln w,\qquad\pi_{HF}=\left(\frac{b_{HF}/A_{F}}{P_{H}}\right)^{-\theta}=\frac{(w/b_{HF})^{\theta}}{(w/b_{HF})^{\theta}+A^{\theta}}=1-\pi_{HH},$ $\mathclap{\forall j}$
@@
$d\ln\omega_{n}^{j}=(1-\sigma)\left(d\ln P_{n}^{j}-\sum_{h=1}^{J}\omega_{n}^{h}d\ln P_{n}^{h}\right),\qquad\omega_{n}^{j}=\frac{(P_{n}^{j})^{1-\sigma}}{\sum_{h=1}^{J}(P_{n}^{h})^{1-\sigma}}=\frac{1}{J},$ $\mathclap{\forall(n,j)}$
@@ Appendix B Proof of Proposition 2
$=\frac{(1-\sigma)\xi(x)}{\sum_{j=1}^{k}\omega^{j}}\sum_{h=1}^{k}\omega^{h}\sum_{j=k+1}^{J}\left(\epsilon^{h}-\epsilon^{j}\right)\omega^{j},\qquad k\in\{1,\dots,J-1\}$
@@
$d\ln P_{F}^{j}=\pi_{FH}d\ln w,\qquad\pi_{FH}=\left(\frac{wb_{FH}/A_{H}}{P_{F}}\right)^{-\theta}=\frac{(A/b_{FH})^{\theta}}{w^{\theta}+(A/b_{FH})^{\theta}}=1-\pi_{FF},$ $\mathclap{\forall j}$
$d\ln P_{H}^{j}=\pi_{HF}d\ln b_{HF}^{j}+\pi_{HH}d\ln w,\qquad\pi_{HF}=\left(\frac{b_{HF}/A_{F}}{P_{H}}\right)^{-\theta}=\frac{(w/b_{HF})^{\theta}}{(w/b_{HF})^{\theta}+A^{\theta}}=1-\pi_{HH},$ $\mathclap{\forall j}$
@@
$d\ln\omega_{n}^{j}=(1-\sigma)\left(d\ln P_{n}^{j}-\sum_{h=1}^{J}\omega_{n}^{h}d\ln P_{n}^{h}\right),\qquad\omega_{n}^{j}=\frac{(P_{n}^{j})^{1-\sigma}}{\sum_{h=1}^{J}(P_{n}^{h})^{1-\sigma}}=\frac{1}{J},$ $\mathclap{\forall(n,j)}$
@@ Appendix C Proof of Proposition 3
$\text{(c)}\ -\frac{1}{va_{H}^{j}}d\left(\frac{\widetilde{T}_{H}^{j}}{w_{H}L_{H}}\right)=-\pi_{HF}d\tau_{HF}<0,\qquad\qquad\qquad\qquad\forall j$
@@
$\text{(c)}\ -\frac{1}{va_{F}^{j}}d\left(\frac{\widetilde{T}_{F}^{j}}{w_{F}L_{F}}\right)=0,\qquad\qquad\qquad\qquad j\in\{1,\dots,J\}$
@@
$\text{(a)}\ \frac{P_{F}^{j}C_{F}^{j}}{w_{F}L_{F}^{j}}d\ln\omega_{F}^{j}=0,\qquad\text{(b)}\ -\frac{1}{va_{F}^{j}}d\left(\frac{\widetilde{T}_{F}^{j}}{w_{F}L_{F}}\right)=0,\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\forall j$
@@ Appendix I Applying the Newton Method
The dynamic welfare implication for the U.S. is positive but smaller in magnitude compared to the permanent shock (0.06 percent). This is simply because the time period during which the U.S. receives higher tariff revenue is shorter in the case of the temporary tariff shock. Canada, India, and Mexico remain the three worst-off countries, but under the temporary shock, welfare loss is largest for Mexico (0.27 percent), followed by India (0.23 percent) and Canada (0.19 percent).

### v4

@@ 5.2 Trade War
Next, we consider a trade war scenario: the U.S. imposes an additional 20-percentage-point tariff on all the manufacturing imports, and every other country retaliates against the U.S. by imposing tariffs on U.S. exports with the same magnitude. As before, the trade war is assumed to come as a surprise shock. Results are displayed in Figure A8 in Supplemental Appendix L. Even with retaliatory tariffs from around the world, U.S. real per capita consumption rises for the first 11 years relative to the baseline, but then turns negative. The saving rate is lower than the baseline equilibrium, which follows the same intuition as in the unilateral tariffs. Implications for sectoral expenditure shares and value-added shares remain the same qualitatively, but are smaller in magnitude. The manufacturing value-added share rises by 0.7–0.9 percentage point relative to the baseline, as compared to 0.9–1.3 percentage points in the case of the unilateral tariff. As the retaliatory scenario is more plausible in the real world, our result suggests that the sectoral implications of tariffs on manufacturing will be more marginal in a quantitative sense.
Welfare implications of the trade war are summarized in the fourth and fifth columns of Table 1. We confirm that under the trade war scenario, all countries are worse off compared to the baseline equilibrium. The U.S. welfare loss is 0.14 percent, which is larger than the welfare losses of most European countries and Japan. Canada has the largest welfare loss of 1.21 percent.
@@ 6 Conclusion
The model is used to simulate a 20-percentage-point increase in U.S. manufacturing tariffs on all trading partners beginning in 2001. We find that the relative price effect and the trade protection effect dominate the income effect. Specifically, U.S. manufacturing value-added share rises by 0.9–1.3 percentage points, while the service sector declines by 0.8–1.1 percentage point between 2001 and 2014. Despite these shifts supporting protectionist objectives, the welfare impact is quantitatively minor. The U.S. is better off only by 0.41 percent at the expense of all other countries. In particular, Canada’s welfare falls by 1.44 percent. If the other countries retaliate with 20 percentage-point higher tariffs, all countries are worse off, with the U.S. welfare loss being 0.14 percent.
@@
Hayato Kato [^1] Graduate School of Economics, the University of Osaka; CESifo Research Network. Email: hayato.kato@econ.osaka-u.ac.jp     Kensuke Suzuki [^2] Department of Economics, Clark University; Economic Research Center, Nagoya University. Email: ksuzuki@clarku.edu     Motoaki Takahashi [^3] Graduate School of Economics, the University of Osaka. Email: motoaki.takahashi@econ.osaka-u.ac.jp
@@ Appendix A Proof of Proposition 1
$\hskip 18.49988pt\hskip 18.49988pt\frac{\theta}{1+\mu_{H}}\frac{d\ln\mu_{H}}{d\tau_{HF}}-\frac{\theta\pi_{HF}w^{\theta}\left[(1+\theta)(w/b_{HF})^{\theta}+A^{\theta}\left(\theta(1+\tau_{HF})+(b_{HF}b_{FH})^{-\theta}\right)\right]}{\Gamma(1+\tau_{HF})},$
@@
$\hskip 18.49988pt\hskip 18.49988pt\hskip 18.49988pt\qquad\ \ =\begin{cases}\geq 0&\text{if}\ \tau_{HF}\leq(\theta\pi_{FF})^{-1}+\Gamma^{\prime}\ \text{or}\ \tau_{HF}\leq\tau_{HF}^{**}\\ <0&\text{if}\ \tau_{HF}>(\theta\pi_{FF})^{-1}+\Gamma^{\prime}\ \text{or}\ \tau_{HF}>\tau_{HF}^{**}\end{cases},$
@@
$d\ln P_{F}^{j}=\pi_{FH}d\ln w,\hskip 18.49988pt\pi_{FH}=\left(\frac{wb_{FH}/A_{H}}{P_{F}}\right)^{-\theta}=\frac{(A/b_{FH})^{\theta}}{w^{\theta}+(A/b_{FH})^{\theta}}=1-\pi_{FF},$ $\mathclap{\forall j}$
$d\ln P_{H}^{j}=\pi_{HF}d\ln b_{HF}^{j}+\pi_{HH}d\ln w,\hskip 18.49988pt\pi_{HF}=\left(\frac{b_{HF}/A_{F}}{P_{H}}\right)^{-\theta}=\frac{(w/b_{HF})^{\theta}}{(w/b_{HF})^{\theta}+A^{\theta}}=1-\pi_{HH},$ $\mathclap{\forall j}$
@@
$d\ln\omega_{n}^{j}=(1-\sigma)\left(d\ln P_{n}^{j}-\sum_{h=1}^{J}\omega_{n}^{h}d\ln P_{n}^{h}\right),\hskip 18.49988pt\omega_{n}^{j}=\frac{(P_{n}^{j})^{1-\sigma}}{\sum_{h=1}^{J}(P_{n}^{h})^{1-\sigma}}=\frac{1}{J},$ $\mathclap{\forall(n,j)}$
@@ Appendix B Proof of Proposition 2
$=\frac{(1-\sigma)\xi(x)}{\sum_{j=1}^{k}\omega^{j}}\sum_{h=1}^{k}\omega^{h}\sum_{j=k+1}^{J}\left(\epsilon^{h}-\epsilon^{j}\right)\omega^{j},\hskip 18.49988ptk\in\{1,\dots,J-1\}$
@@
$d\ln P_{F}^{j}=\pi_{FH}d\ln w,\hskip 18.49988pt\pi_{FH}=\left(\frac{wb_{FH}/A_{H}}{P_{F}}\right)^{-\theta}=\frac{(A/b_{FH})^{\theta}}{w^{\theta}+(A/b_{FH})^{\theta}}=1-\pi_{FF},$ $\mathclap{\forall j}$
$d\ln P_{H}^{j}=\pi_{HF}d\ln b_{HF}^{j}+\pi_{HH}d\ln w,\hskip 18.49988pt\pi_{HF}=\left(\frac{b_{HF}/A_{F}}{P_{H}}\right)^{-\theta}=\frac{(w/b_{HF})^{\theta}}{(w/b_{HF})^{\theta}+A^{\theta}}=1-\pi_{HH},$ $\mathclap{\forall j}$
@@
$d\ln\omega_{n}^{j}=(1-\sigma)\left(d\ln P_{n}^{j}-\sum_{h=1}^{J}\omega_{n}^{h}d\ln P_{n}^{h}\right),\hskip 18.49988pt\omega_{n}^{j}=\frac{(P_{n}^{j})^{1-\sigma}}{\sum_{h=1}^{J}(P_{n}^{h})^{1-\sigma}}=\frac{1}{J},$ $\mathclap{\forall(n,j)}$
@@ Appendix C Proof of Proposition 3
$\text{(c)}\ -\frac{1}{va_{H}^{j}}d\left(\frac{\widetilde{T}_{H}^{j}}{w_{H}L_{H}}\right)=-\pi_{HF}d\tau_{HF}<0,\hskip 18.49988pt\hskip 18.49988pt\hskip 18.49988pt\hskip 18.49988pt\forall j$
@@
$\text{(c)}\ -\frac{1}{va_{F}^{j}}d\left(\frac{\widetilde{T}_{F}^{j}}{w_{F}L_{F}}\right)=0,\hskip 18.49988pt\hskip 18.49988pt\hskip 18.49988pt\hskip 18.49988ptj\in\{1,\dots,J\}$
@@
$\text{(a)}\ \frac{P_{F}^{j}C_{F}^{j}}{w_{F}L_{F}^{j}}d\ln\omega_{F}^{j}=0,\hskip 18.49988pt\text{(b)}\ -\frac{1}{va_{F}^{j}}d\left(\frac{\widetilde{T}_{F}^{j}}{w_{F}L_{F}}\right)=0,\hskip 18.49988pt\hskip 18.49988pt\hskip 18.49988pt\hskip 18.49988pt\hskip 18.49988pt\hskip 18.49988pt\hskip 18.49988pt\hskip 18.49988pt\forall j$
@@ Appendix I Applying the Newton Method
The dynamic welfare implication for the U.S. is positive but smaller in magnitude compared to the permanent shock (0.06 percent). This is simply because the time period during which the U.S. receives higher tariff revenue is shorter in the case of the temporary tariff shock. Canada, India, and Mexico remain the three worst-off countries, but under the temporary shock, welfare loss is largest for Mexico (0.27 percent), followed by India (0.25 percent) and Canada (0.19 percent).